2. Find the area of the region. One petal of r = 2 cos 30
Answer #1

0=0 8 8 = 2 C3 (30) 8=2 O = + 7/ 8=12 O= + 7 8 = = √2-12 4 0= + 1 - 28=0 - Area of one petal, A = ( 2 1 0 ( do) x 2 (2=0, B =